A beam of particles strikes a wall containing 2x10^29 nuclei per m^3. Each nucleus behaves like a sphere of radius 3x10^−15 m. Find the thickness of a wall that exactly half the particles will penetrate without scattering. What thickness would be needed to stop all but one particle in 10^6 ?


1
Expert's answer
2022-05-12T15:09:06-0400

"d=\frac{\sqrt\frac rn}2=6.1\cdot 10^{-8}~m."